A VW Beetle goes from 0 to 62.8 mi/h with an acceleration of +2.33 m/s2.

(a) How much time does it take for the Beetle to reach this speed?

(b) A top-fuel dragster can go from 0 to 60.0 mi/h in 0.600 s. Find the acceleration of the dragster.
Answer #1

A)here we have v= 62.8mph = 28.074m/s, u =0

v = u +at

28.074 =0 + 2.33t

t = 28.074/2.33

= 12.05 s

B) Again,

a = v/t

= 26.822/0.6

= 44.7 m/s^2